Find the points which trisect the segment joining the points (-3,5) and (9,2).
1
Expert's answer
2011-11-18T08:36:26-0500
We'll use Fales& theorem
(x,y), (a,b)
sqrt((-3-x)^2+(5-y)^2)=sqrt((x-a)^2+(y-b)^2)=sqrt((a-9)^2+(b-2)^2)
(X,y)=(1,4)
(a,b)=(5,3)